OpenStudy (anonymous):

(s+(1)/(4)) (s+(1)/(4)) multiply

OpenStudy (dumbcow):

\[\frac{s+1}{4}*\frac{s+1}{4}=\frac{s^{2}+2s+1}{16}\]